OpenStudy (mckennarian):

CHECK MY WORK? Let f(x) = 4x^2 + x + 1 and g(x) = x2 – 2. Find g(f(x)). Show each step of your work. f(x)= 4x^2 + x + 1 g(x)= x^2 - 2 g(f(x))= (4x^2 + x + 1)^2 - 2 g(f(x))= 16x^4 + x^2 + 1^2 + 3x^2 g(fx))= 16x^4 + 8x^3 + 9x^2 + 2x - 1

OpenStudy (anonymous):

@haley15122

OpenStudy (tyteen4a03):

I'm confused; what exactly is line 4 doing?

OpenStudy (irishboy123):

not sure how you did that but FWIW my working gets you to same place.... \(g(f(x))= (4x^2 + x + 1)^2 - 2\) ASIDE: \((4x^2 + x + 1)^2 = (4x^2 + (x + 1))^2 \) i don't see this in your solution....? \(= 16x^4 + (2)4x^2(x+1)+ (x+1)^2 -2\) \(= 16x^4 + 8x^3 + 8x^2 + x^2 + 2x + 1-2\) \(= 16x^4 + 8x^3 + 9x^2 +2x -1\)